OpenStudy (anonymous):

(sin^2(theta)=cos^2(theta))/cos^2(theta)=sec^2(theta)

OpenStudy (dumbcow):

\[\frac{\sin^{2} \theta - \cos^{2} \theta}{\cos^{2} \theta} = \sec^{2} \theta\]

OpenStudy (dumbcow):

or that should be a "+"